OpenStudy (anonymous):

is 60.796 the correct answer for sqrt3 (7sqrt2-9sqrt25)

OpenStudy (dumbcow):

yes but it should be negative

OpenStudy (anonymous):

what would be the exact answer?

OpenStudy (anonymous):

61?

OpenStudy (dumbcow):

what do you mean? depends on how many decimal places is needed

OpenStudy (anonymous):

I get it now thanks :) Whatever decimal places the facilitator asks for and it would be rounded!

OpenStudy (dumbcow):

right :)

OpenStudy (anonymous):

this whole problem is under the radical sqrt x^2-14x+49

OpenStudy (dumbcow):

factor x^2-14x +49 its a perfect square, sqrt(49) = 7, 2*7 = 14 --> sqrt(x-7)^2) = x-7

OpenStudy (anonymous):

index4 sqrt 10,000*index4 sqrt 100,000